# From Lipschitz Embedding to Lipschitz Equivalence between Dust-like Self-similar Sets

### v1

@@ 1. Introduction
Self-similar sets, which arise as attractors of self-similar iterated function systems (IFS), exhibit intricate geometric and topological properties that make them fundamental objects in the study of fractals. A key question in this context is to complete the Lipschitz classification of these sets, as Lipschitz equivalence preserves a number of crucial geometric and measure-theoretic properties, including fractal dimensions, regularity and metric scaling structures. Recall that $A,B\subset\mathbb{R}^{d}$ are said to be Lipschitz equivalent if there exists a bijection $f:A\to B$ such that
@@
Unlike topological equivalence, the Lipschitz equivalence requires a bounded geometric distortion, making the classification far more rigid and heavily reliant on the precise local patterns of self-similar sets. A natural starting point is to examine dust-like self-similar sets, where the absence of overlaps yields a clear geometric structure. Recall that a self-similar set $K$ generated by an iterated function system (IFS) $\Phi=\{\varphi_{i}\}_{i=1}^{n}$, where each $\varphi_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a contractive similitude, is called dust-like when the images $\{\varphi_{i}(K)\}_{i=1}^{n}$ are pairwise disjoint. In this case, we also say that $\Phi$ satisfies the strong separation condition (SSC). For dust-like self-similar sets, it is proved that the Lipschitz equivalence reduces entirely to the algebraic compatibility of scaling ratios in their associated IFSs (see e.g. [15]), thus simplifying many technical challenges. The situation becomes particularly tractable ((see Lemma 2.4)) when dealing with homogeneous self-similar sets, where all similitudes in the associated IFS share a common contraction ratio.
@@
One of the foundational studies toward this algebraic direction was conducted by Falconer and Marsh [8]. They demonstrated that the Lipschitz equivalence between two dust-like self-similar sets requires specific multiplicative relationships among the corresponding scaling ratios, thereby revealing a deep connection to the algebraic properties of the IFSs. This line was further developed in [14] where the first author, in collaboration with Rao and Wang, introduced a necessary condition for the Lipschitz equivalence called matchable condition, which enables a complete classification for two important cases: when the contraction vectors have full ranks, or when the IFSs consist of exactly two maps. Later, Rao and Zhang [17] linked the classification with a higher-dimensional Frobenius problem, yielding criteria for equivalence when the scaling ratios are coplanar in a certain sense. More recently, Xi and Xiong [22] gave a checkable characterization for systems with commensurable ratios. For related results, see [3, 4, 12, 20] and the references therein.
@@
Another direction investigates whether a dust-like self-similar set admits a Lipschitz embedding into another, and how such embeddability relates to the Lipschitz equivalence. Since Lipschitz maps cannot increase Hausdorff dimension, if $K,F$ are two dust-like self-similar sets with $\operatorname{dim_{H}}K<\operatorname{dim_{H}}F$ (where $\operatorname{dim_{H}}$ denotes the Hausdorff dimension), then no Lipschitz maps can send $K$ onto $F$. On the other hand, if $\operatorname{dim_{H}}K>\operatorname{dim_{H}}F$, a nontrivial Lipschitz embedding of $K$ into $F$ always exists (see [2]). The critical case occurs when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$. For this scenario, Deng, Wen, Xiong and Xi [5] proved that $K$ admits a bi-Lipschitz embedding into $F$ if and only if these sets are Lipschitz equivalent.
@@
The construction of the desired surjection proceeds via three steps. First, we identify a regular local region within the intersection $f(K)\cap F$ that is not too condensed in an appropriate sense. Next, we apply a rescaling argument based on a standard density theorem in geometric measure theory. Finally, we establish the existence of the required map by passing to an appropriate limit. Combining this result with Theorem 1.1, we immediately obtain an affirmative answer to Question 1.2(1).
@@
Let $K,F$ be as in Question 1.2. If $K$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K,F$ are Lipschitz equivalent.
Regarding the second question, our findings present a dichotomy of results. On the one hand, we prove that the Lipschitz equivalence holds when $F$ (as the range of the Lipschitz embedding) is homogeneous, establishing a converse to Balka and Keleti’s original theorem. On the other hand, we demonstrate that the homogeneity assumption on $K$ or $F$ is indeed essential: without it, there exists an explicit example (detailed in the last section) that provides a negative answer to Question 1.2(2). We summarize these results as follows.
@@
    If $F$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K,F$ are Lipschitz equivalent.
@@
    Assume that $K$ can be sent onto $F$ by a Lipschitz map. Without the homogeneity assumption, $K,F$ need not to be Lipschitz equivalent.
Here is an outline of the key steps of the proof. For part (1), it suffices to build the algebraic dependence between the corresponding contraction ratios (see Lemma 2.4). Note that Theorem 1.3 guarantees a Lipschitz surjection from $K$ to $F$. From this, we construct an alternative map which is almost injective and enjoys a crucial measure-preserving property. It is worth pointing out that the measure-preserving property is enough to derive the algebraic dependence when $K$ is homogeneous, thus providing an alternative way to recover Theorem 1.1. When $F$ is homogeneous instead, the argument admits a certain “reversal” (in some sense) once we prove that the image of every cell of $K$ under the alternative map contains a cell of $F$ of comparable size. For part (2), we rely on two powerful tools from [22]: a criterion of the equivalence in the commensurable case and a mass decomposition lemma. Using these, we explicitly construct a Lipschitz surjection between two dust-like self-similar sets that fail to be Lipschitz equivalent, thereby completing the proof.
Our results are also related to a notable conjecture by Feng, Huang and Rao [9], which states that an affine embedding from $K$ into $F$ may enforce the algebraic dependence between their associated ratios even without assuming that $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$. Significant progress has been made toward this conjecture, see [1, 6, 10]. In particular, a resolution of the special case when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$ was obtained as a corollary in an ongoing work [16] of the second author joint with Rao. The proof of Theorems 1.1, 1.3 and 1.5(1) reveals that if in addition that either $K$ or $F$ is homogeneous, the algebraic dependence follows already from the existence of a nontrivial Lipschitz embedding—a weaker condition than the original affinity requirement. However, as shown in [5], the ratios may remain independent without the dimensional restriction.
The paper is organized as follows. Section 2 introduces some necessary preliminaries. Section 3 establishes our notational conventions and presents the proof of Theorem 1.3. Building on these foundations, Sections 4 and 5 develop the construction of a well-behaved Lipschitz surjection from $K$ to $F$ and present the proof of Theorem 1.5(1). Finally, Section 6 provides an explicit example that demonstrates the second part of Theorem 1.5.
@@ 2. Preliminaries
In this section we prove some useful preliminaries. We start with recalling the classical Lipschitz extension theorem of Kirszbraun.
@@
For any sequence $\{K_{n}\}_{n=1}^{\infty}\subset\mathcal{K}$ and Lipschitz maps $f_{n}:K_{n}\to F$ with $\sup_{n}\operatorname{Lip}(f_{n})\leq L$, there exists $E\in\mathcal{K}$ and a Lipschitz map $g:E\to F$ with $\operatorname{Lip}(g)\leq L$ such that
$\mathcal{H}^{s}(E)=\lim_{j\to\infty}\mathcal{H}^{s}(K_{n_{j}})\quad\text{and}\quad\mathcal{H}^{s}(g(E))\geq\limsup_{j\to\infty}\mathcal{H}^{s}(f_{n_{j}}(K_{n_{j}}))$
@@
Since $K_{n}\in\mathcal{K}$, we can write it as $K_{n}=\bigcup_{t=1}^{p_{n}}(c_{n,t}O_{n,t}K+a_{n,t})$. Passing to a subsequence if necessary, we may assume that $p_{n}\equiv p$ for some $1\leq p\leq P$. Put $\widetilde{\gamma}:=v+L^{-1}+\gamma$ and $b_{t}:=(3t\widetilde{\gamma},0,\ldots,0)\in\mathbb{R}^{d}$ for $1\leq t\leq p$. For $n\geq 1$, it is easy to see that when $t\neq t^{\prime}$,
@@
$\mathcal{H}^{s}(E)=\sum_{t=1}^{p}c_{t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\sum_{t=1}^{p}c_{n,t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\mathcal{H}^{s}(K_{n}).$
To find a limit of $g_{n}$, we first pick a large $R>0$ so that $\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\subset B(0,R)$ for all $n\geq 1$. By Lemma 2.1, we can extend each $g_{n}$ to a Lipschitz map $\widetilde{g}_{n}:B(0,R)\to\mathbb{R}^{d}$ with $\operatorname{Lip}(\widetilde{g}_{n})=\operatorname{Lip}(g_{n})\leq L$. In particular, $\{\widetilde{g}_{n}\}$ is equicontinuous. Furthermore, since $\widetilde{g}_{n}(B(0,R))\cap F\neq\varnothing$ and
@@
$\widetilde{g}_{n}(B(0,R))$ is contained in the $2LR$-neighborhood of $F$ for all $n$, that is, $\{\widetilde{g}_{n}\}$ is pointwise bounded. Then by Arzelá-Ascoli theorem (e.g., see [19, Corollary 1.8.25]), there exists a subsequence $\{\widetilde{g}_{n_{j}}\}_{j=1}^{\infty}$ which converges uniformly to a limit function, say $g$. Clearly, $g$ is also Lipschitz with $\operatorname{Lip}(g)\leq L$.
@@
where $\mathcal{N}_{\varepsilon}(\cdot)$ denotes the open $\varepsilon$-neighborhood. Since $\mathcal{H}^{s}(F)<\infty$, we see that
(2.4) $\mathcal{H}^{s}(g(E))=\lim_{\varepsilon\to 0}\mathcal{H}^{s}(\mathcal{N}_{\varepsilon}(g(E))\cap F)$
and thus
@@
Write $\lambda_{j}=r^{\alpha_{j}}$, where $\alpha_{j}\in\mathbb{Q}$ for $j\in J$. The SSC implies that $I\cdot r^{s}=1$ and $\sum_{j\in J}\lambda_{j}^{s}=1$. If $I=m^{k}$ for some integers $m,k>1$, then $K$ is Lipschitz equivalent to any self-similar set $E=\bigcup_{t=1}^{m}g_{t}(E)$ with the SSC such that each $g_{t}$ has similarity ratio $r^{1/k}$ (e.g., see [15]). So we may assume that $k=1$. Then $r^{s}=m^{-1}$ and thus
@@
It follows from [2, Lemma 7.1] that $\alpha_{j}$ are all integers. Then by [2, Theorem 7.7], $K$ and $F$ are Lipschitz equivalent. ∎
@@
  • $I^{*}:=\bigcup_{n=1}^{\infty}I^{n}$ and $K_{\bm{i}}:=\varphi_{\bm{i}}(K)$ for $\bm{i}\in I^{*}$;
  • $J^{*}:=\bigcup_{n=1}^{\infty}J^{n}$ and $F_{\bm{j}}:=\psi_{\bm{j}}(F)$ for $\bm{j}\in J^{*}$;
  • For $\bm{i}=i_{1}\cdots i_{n}\in I^{*}$, $\varphi_{\bm{i}}:=\varphi_{i_{1}}\circ\cdots\circ\varphi_{i_{n}}$ and $r_{\bm{i}}=r_{i_{1}}\cdots r_{i_{n}}$;
  • For $\bm{j}=j_{1}\cdots j_{n}\in J^{*}$, $\psi_{\bm{j}}:=\psi_{j_{1}}\circ\cdots\circ\psi_{j_{n}}$ and $\lambda_{\bm{j}}:=\lambda_{j_{1}}\cdots\lambda_{j_{n}}$;
  • $\underline{r}:=\min_{i}r_{i}$ and $\underline{\lambda}:=\min_{j}\lambda_{j}$;
  • $\wedge$: the operation between any two words in $I^{*}$ or $J^{*}$ which produces their longest common prefix;
  • For $\bm{i}\in I^{*}$ (resp. $\bm{j}\in J^{*}$), write $\bm{i}^{-}$ (resp. $\bm{j}^{-}$) to be the word obtained by removing the last digit of $\bm{i}$ (resp. $\bm{j}$);
  • $|\cdot|$: either the modulus of a vector in $\mathbb{R}^{d}$ or the length of a finite word in $I^{*}\cup J^{*}$;
  • $\Delta_{K}:=\min_{i\neq i^{\prime}}\operatorname{dist}(K_{i},K_{i^{\prime}})$ and $\Delta_{F}:=\min_{j\neq j^{\prime}}\operatorname{dist}(F_{j},F_{j^{\prime}})$, which are both positive numbers due to the SSC.
@@
We begin with constructing a map $g:K\to F$ with $g(K)=f(K)\cap F$. For any $\bm{i}\in I^{*}$ such that $f(K_{\bm{i}})\cap F=\varnothing$ but $f(K_{\bm{i}^{-}})\cap F\neq\varnothing$, we choose an arbitrary point $y_{\bm{i}}$ in $f(K_{\bm{i}^{-}})\cap F$ and define $g(K_{\bm{i}})=\{y_{\bm{i}}\}$. So if $x\in K$ is such that $f(x)\notin F$, then $g(x)$ is defined as $y_{\omega}$, where $\omega$ is the unique word in $I^{*}$ such that $x\in K_{\omega}$, $f(K_{\omega})\cap F=\varnothing$ and $f(K_{\omega^{-}})\cap F\neq\varnothing$. Note that the existence of $\omega$ is due to the compactness of $F$. If $f(x)\in F$ we simply set $g(x)=f(x)$. In this way, the map $g:K\to F$ is well defined and it is clear that $g(K)=f(K)\cap F$.
We claim that $\operatorname{diam}(g(K_{\bm{i}}))\leq\operatorname{diam}(f(K_{\bm{i}}))$ for all $\bm{i}\in I^{*}$. In fact, if $f(K_{\bm{i}})\cap F=\varnothing$, then $g(K_{\bm{i}})$ is a singleton and the inequality clearly holds. If $f(K_{\bm{i}})\cap F\neq\varnothing$, then by the definition of $g$,
@@
It remains to prove that $g$ is Lipschitz. For any distinct $x,x^{\prime}\in K$, let $\omega\in I^{*}$ be such that $|\omega|=\max\{|\bm{i}|:\bm{i}\in I^{*},x,x^{\prime}\in K_{\bm{i}}\}$. So
@@
Based on this, the proof of Theorem 1.3 reduces to establishing the following result.

### v2

@@ 1. Introduction
Self-similar sets, which arise as attractors of self-similar iterated function systems (IFS), exhibit intricate geometric and topological properties that make them fundamental objects in the study of fractals. A key question in this context is to complete the Lipschitz classification of such sets, as Lipschitz equivalence preserves a number of crucial geometric and measure-theoretic properties, including fractal dimensions, regularity and metric scaling structures. Recall that $A,B\subset\mathbb{R}^{d}$ are said to be Lipschitz equivalent if there exists a bijection $f:A\to B$ such that
@@
Unlike topological equivalence, the Lipschitz equivalence requires a bounded geometric distortion, making the classification far more rigid and heavily reliant on the precise local patterns of self-similar sets. A natural starting point is to examine dust-like self-similar sets, where the absence of overlaps yields a clear geometric structure. Recall that a self-similar set $K$ generated by an iterated function system (IFS) $\Phi=\{\varphi_{i}\}_{i=1}^{n}$, where each $\varphi_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a contractive similitude, is called dust-like when $n\geq 2$ and the images $\{\varphi_{i}(K)\}_{i=1}^{n}$ are pairwise disjoint. In this case, we also say that $\Phi$ satisfies the strong separation condition (SSC). For dust-like self-similar sets, it is proved that the Lipschitz equivalence reduces entirely to the algebraic compatibility of scaling ratios in their associated IFSs (see e.g. [15]), thus simplifying many technical challenges. The situation becomes particularly tractable (see Lemma 2.4) when dealing with homogeneous self-similar sets, where all similitudes in the associated IFS share a common contraction ratio.
@@
One of the foundational studies toward the algebraic direction was conducted by Falconer and Marsh [8]. They demonstrated that the Lipschitz equivalence between two dust-like self-similar sets requires specific multiplicative relationships among the corresponding scaling ratios, thereby revealing a deep connection to the algebraic properties of the IFSs. This line was further developed in [14] where the first author, in collaboration with Rao and Wang, introduced a necessary condition for the Lipschitz equivalence called matchable condition, which enables a complete classification for two important cases: when the contraction vectors have full ranks, or when the IFSs consist of exactly two maps. Later, Rao and Zhang [17] linked the classification with a higher-dimensional Frobenius problem, yielding criteria for equivalence when the scaling ratios are coplanar in a certain sense. More recently, Xi and Xiong [22] gave a checkable characterization for systems with commensurable ratios. For related results, see [3, 4, 12, 20] and the references therein.
@@
Another direction investigates Lipschitz embeddings between dust-like self-similar sets and how such embeddability relates to the Lipschitz equivalence. Since Lipschitz maps cannot increase Hausdorff dimension, if $K,F$ are two dust-like self-similar sets with $\operatorname{dim_{H}}K<\operatorname{dim_{H}}F$ (where $\operatorname{dim_{H}}$ denotes the Hausdorff dimension), then no Lipschitz maps can send $K$ onto $F$. On the other hand, if $\operatorname{dim_{H}}K>\operatorname{dim_{H}}F$, a Lipschitz embedding of $K$ onto $F$ always exists (see [2]). The critical case occurs when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$. For this scenario, Deng, Wen, Xiong and Xi [5] proved that $K$ admits a bi-Lipschitz embedding into $F$ if and only if $K$ and $F$ are Lipschitz equivalent.
@@
The construction of such a surjection proceeds via three steps. We begin by determining a regular local region of $F$ where the map $f$ is not overly concentrated. Next, we apply a rescaling argument based on a standard density theorem in geometric measure theory. The desired surjection is then constructed by passing to an appropriate limit. Combining this result with Theorem 1.1, we immediately obtain an affirmative answer to Question 1.2(1).
@@
Let $K,F$ be as in Question 1.2. If $K$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K$ and $F$ are Lipschitz equivalent.
Regarding the second question, our findings present a dichotomy of results. On the one hand, we prove that the Lipschitz equivalence holds when $F$ (as the target space of the embedding) is homogeneous, establishing a converse to Balka and Keleti’s original theorem. On the other hand, we demonstrate that the homogeneity assumption on $K$ or $F$ is indeed essential: without it, there exists an explicit example (see Example 6.2) that provides a negative answer to Question 1.2(2). We summarize these results as follows.
@@
    If $F$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K$ and $F$ are Lipschitz equivalent.
@@
    Assume that $K$ can be sent onto $F$ by a Lipschitz map. Without the homogeneity assumption, $K$ and $F$ need not to be Lipschitz equivalent.
Here is an outline of the key steps of the proof. For part (1), it suffices to build the algebraic dependence between the corresponding contraction ratios (see Lemma 2.4). Note that Theorem 1.3 guarantees a Lipschitz surjection from $K$ to $F$. From this, we construct an alternative map which is almost injective and enjoys a crucial measure-preserving property. It is worth pointing out that the measure-preserving property is enough to derive the algebraic dependence when $K$ is homogeneous, thus providing an alternative way to recover Theorem 1.1. When $F$ is homogeneous instead, the argument admits a certain “reversal” (in some sense) once we prove that the image of every cell of $K$ under the alternative map contains a cell of $F$ of comparable size. For part (2), we rely on a powerful mass decomposition lemma in [22]. Using these, we explicitly construct a Lipschitz surjection between two dust-like self-similar sets that fail to be Lipschitz equivalent, thereby completing the proof.
Our results are also related to a notable conjecture by Feng, Huang and Rao [9], which states that the existence of an affine embedding from $K$ into $F$ may enforce the algebraic dependence between their associated ratios even without assuming that $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$. Significant progress has been made toward this conjecture, see [1, 6, 10]. In particular, a resolution of the special case when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$ was recently obtained as a corollary in an ongoing work [16] of the second author joint with Rao. The proofs of Theorems 1.1, 1.3 and 1.5(1) reveal that if in addition that either $K$ or $F$ is homogeneous, the algebraic dependence follows already from the existence of a nontrivial Lipschitz embedding—a weaker condition than the affinity requirement. However, as shown in [5], such dependence is not guaranteed without the dimensional restriction.
The paper is organized as follows. Section 2 introduces some necessary preliminaries. Section 3 establishes our notational conventions and presents the proof of Theorem 1.3. Building on these foundations, Sections 4 and 5 develop the construction of a well-behaved Lipschitz surjection and present the proof of Theorem 1.5(1). Finally, Section 6 provides an explicit example that demonstrates the second part of Theorem 1.5.
@@ 2. Preliminaries
We begin with a classical Lipschitz extension theorem of Kirszbraun.
@@
For any sequence $\{E_{n}\}_{n=1}^{\infty}\subset\mathcal{K}$ and Lipschitz maps $f_{n}:E_{n}\to F$ with $\sup_{n}\operatorname{Lip}(f_{n})\leq L$, there exists $E\in\mathcal{K}$ and a Lipschitz map $g:E\to F$ with $\operatorname{Lip}(g)\leq L$ such that
$\mathcal{H}^{s}(E)=\lim_{j\to\infty}\mathcal{H}^{s}(E_{n_{j}})\quad\text{and}\quad\mathcal{H}^{s}(g(E))\geq\limsup_{j\to\infty}\mathcal{H}^{s}(f_{n_{j}}(E_{n_{j}}))$
@@
Since $E_{n}\in\mathcal{K}$, we can write it as $E_{n}=\bigcup_{t=1}^{p_{n}}(c_{n,t}O_{n,t}K+a_{n,t})$. Passing to a subsequence if necessary, we may assume that $p_{n}\equiv p$ for some $1\leq p\leq P$. Put $\widetilde{\gamma}:=v+L^{-1}+\gamma$ and $b_{t}:=(3t\widetilde{\gamma},0,\ldots,0)\in\mathbb{R}^{d}$ for $1\leq t\leq p$. For $n\geq 1$, it is easy to see that when $t\neq t^{\prime}$,
@@
$\mathcal{H}^{s}(E)=\sum_{t=1}^{p}c_{t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\sum_{t=1}^{p}c_{n,t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\mathcal{H}^{s}(E_{n}).$
To find a limit of $g_{n}$, we first pick a large $R>0$ so that $\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\subset B(0,R)$ for all $n\geq 1$. By Lemma 2.1, each $g_{n}$ extends to a Lipschitz map $\widetilde{g}_{n}:B(0,R)\to\mathbb{R}^{d}$ with $\operatorname{Lip}(\widetilde{g}_{n})=\operatorname{Lip}(g_{n})\leq L$. In particular, $\{\widetilde{g}_{n}\}$ is equicontinuous. Furthermore, since $\widetilde{g}_{n}(B(0,R))\cap F\neq\varnothing$ and
@@
$\widetilde{g}_{n}(B(0,R))$ is contained in the $2LR$-neighborhood of $F$ for all $n$, that is, $\{\widetilde{g}_{n}\}$ is pointwise bounded. Then by Arzelá-Ascoli theorem (e.g., see [19, Corollary 1.8.25]), there exists a subsequence $\{\widetilde{g}_{n_{j}}\}_{j=1}^{\infty}$ which converges uniformly to a limit map, say $g$. Clearly, $g$ is also Lipschitz with $\operatorname{Lip}(g)\leq L$.
@@
where $\mathcal{N}_{\varepsilon}(\cdot)$ denotes the open $\varepsilon$-neighborhood. Note that the restriction of $\mathcal{H}^{s}$ on $F$ is a Radon measure and, since $g(E)\subset F$,
$\mathcal{H}^{s}(g(E))=\mathcal{H}^{s}(g(E)\cap F)=\lim_{\varepsilon\to 0}\mathcal{H}^{s}(\mathcal{N}_{\varepsilon}(g(E))\cap F).$
Therefore,
@@
Write $\lambda_{j}=r^{\alpha_{j}}$, where $\alpha_{j}\in\mathbb{Q}$ for $j\in J$. The SSC implies that $I\cdot r^{s}=1$ and $\sum_{j\in J}\lambda_{j}^{s}=1$. If $I=m^{k}$ for some integers $m,k>1$, then $K$ is Lipschitz equivalent to any self-similar set $E=\bigcup_{t=1}^{m}g_{t}(E)$ with the SSC such that each $g_{t}$ has similarity ratio $r^{1/k}$ (e.g., see [14]). So we may assume that $k=1$ and $m$ is not of the form $p^{\ell}$ where $p,\ell$ are integers strictly larger than $1$. Then $r^{s}=m^{-1}$ and thus
@@
It follows from [2, Lemma 7.1] that $\alpha_{j}$ are all integers. Then by [2, Theorem 7.7] or [14, Theorem 1.5], $K$ and $F$ are Lipschitz equivalent. ∎
@@
Let $I^{*}:=\bigcup_{n=1}^{\infty}I^{n}$ be the set of all finite words over the alphabet $I$ with the convention that $I^{0}=\{\vartheta\}$, where $\vartheta$ denotes the empty word. For $n\geq 0$ and $\bm{i}\in I^{n}$, we call $n$ the length of $\bm{i}$ and denote it by $|\bm{i}|$. When $n\geq 1$, we write for $\bm{i}=i_{1}\cdots i_{n}\in I^{n}$ that
$\varphi_{\bm{i}}:=\varphi_{i_{1}}\circ\cdots\circ\varphi_{i_{n}},\quad K_{\bm{i}}:=\varphi_{\bm{i}}(K),\quad r_{\bm{i}}:=r_{i_{1}}\cdots r_{i_{n}},\quad\bm{i}^{-}:=i_{1}\cdots i_{n-1},$
with the convention that $\bm{i}^{-}=\vartheta$ if $n=1$; when $n=0$, we write $\varphi_{\vartheta}$ to be the identity map and let $K_{\vartheta}:=K$ and $r_{\vartheta}:=1$. For any $\bm{i},\bm{i}^{\prime}\in I^{*}$, $\bm{i}\wedge\bm{i}^{\prime}$ denotes the longest common prefix of $\bm{i}$ and $\bm{i}^{\prime}$. Analogously, for the alphabet $J$, we write $J^{*}:=\bigcup_{n=1}^{\infty}J^{n}$, $J^{0}:=\{\vartheta\}$ and adopt the notation $\psi_{\bm{j}}$, $F_{\bm{j}}$, $\lambda_{\bm{j}}$, $\bm{j}^{-}$ and $\bm{j}\wedge\bm{j}^{\prime}$ in the same manner as above.
Furthermore, let $\underline{r}:=\min_{i\in I}r_{i}$, $\underline{\lambda}:=\min_{j\in J}\lambda_{j}$, $\Delta_{K}:=\min_{i\neq i^{\prime}\in I}\operatorname{dist}(K_{i},K_{i^{\prime}})$ and $\Delta_{F}:=\min_{j\neq j^{\prime}\in J}\operatorname{dist}(F_{j},F_{j^{\prime}})$. The SSC ensures that $\Delta_{K}$ and $\Delta_{F}$ are both positive.
@@
We begin by constructing a map $g:K\to F$ with $g(K)=f(K)\cap F$. For any $\bm{i}\in I^{*}$ such that $f(K_{\bm{i}})\cap F=\varnothing$ but $f(K_{\bm{i}^{-}})\cap F\neq\varnothing$, choose an arbitrary point $y_{\bm{i}}$ in $f(K_{\bm{i}^{-}})\cap F$ and define $g(K_{\bm{i}})=\{y_{\bm{i}}\}$. So for those $x\in K$ with $f(x)\notin F$, $g(x)$ is defined as $y_{\omega}$, where $\omega\in I^{*}$ is the unique word such that $x\in K_{\omega}$, $f(K_{\omega})\cap F=\varnothing$ and $f(K_{\omega^{-}})\cap F\neq\varnothing$. Note that the existence of such an $\omega$ is guaranteed by the closedness of $F$. For $x\in K$ with $f(x)\in F$, we simply set $g(x)=f(x)$. In this way, the map $g:K\to F$ is well defined and it is clear that $g(K)=f(K)\cap F$.
We claim that $\operatorname{diam}(g(K_{\bm{i}}))\leq\operatorname{diam}(f(K_{\bm{i}}))$ for all $\bm{i}\in I^{*}$. Indeed, if $f(K_{\bm{i}})\cap F=\varnothing$, then $g(K_{\bm{i}})$ is a singleton and the inequality holds trivially. If $f(K_{\bm{i}})\cap F\neq\varnothing$, then by the definition of $g$,
@@
It remains to prove that $g$ is Lipschitz. For any distinct $x,x^{\prime}\in K$, let $\omega\in I^{*}\cup I^{0}$ be the longest word such that $x,x^{\prime}\in K_{\omega}$. So
@@
The proof of Theorem 1.3 then reduces to establishing the next result.